How much water must you add to 575.0 mL of orange juice concentrate if you want the final juice to be 15.0% of the strength of the orginal? What is the dilution factor?

*Please show work*
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 100 %

M2 = 15 %

V1 = 575 mL

use:

M1*V1 = M2*V2

V2 = (M1 * V1) / M2

V2 = (100*575)/15

V2 = 3833 mL

Volume of water added = V2 - V1

= 3833 mL - 575 mL

= 3258 mL

Dilution factor = V2/V1

= 3833 / 575

= 6.66

Answer:

Volume of water added = 3258 mL

Dilution factor = 6.66